Express the concentration of a solution (in %w/w, %v/v, and %w/v) containing 5 grams of NaCl and 50 ml of water. (p of NaCl= 2.16 g/ml) p=density.
1
Expert's answer
2020-10-16T08:58:42-0400


Weight of solute (Nacl)=5g

From questions, the volume of water is=50ml

We know that the density of water=1g/cm3

So mass of water=(density)(volume)

=(1gm/cm3)(50ml)

=50g

So,. %(w/w)=(weight of solute)100/(weight of solutions)

%(w/w)=5g(100)/50g

=10

Hence,10%(w/w) Nacl solution.

And also findout the volume of solute(Nacl)

So volume of Nacl=Mass/Density

=5g/(2.16 g/ml)

=2.31ml

And,%(v/v)=(volume of solute)100/(volume of solution)

=(2.31ml)100/50 ml

=4.62


Hence ,4.62%(v/v) Nacl solution.

And ,%(w/v)=(weight of solute)100/volume of solution.

=(5g)100/50ml

=10

Hence, 10%(w/v) Nacl solution.